# Testing the Homogeneity of Two Proportions for Correlated Bilateral Data via the Clayton Copula

### v1

@@ 2.1 Copula Model
Without loss of generality, we use the operation $\overset{*}{=}$ to emphasize the simplification for $\pi_{i1}=\pi_{i2}=\pi_{i}$. This assumption will be hold in the following discussion.
@@ 2.2 Likelihood Inference
$\ell(\pi_{1},\dots,\pi_{g},\theta\mid\tilde{M})$ $=\sum_{i=1}^{g}\Bigg{[}m_{i0}\log\left(C_{\theta}(1-\pi_{i},1-\pi_{i})\right)$
@@
$\ell(\pi_{1},\dots,\pi_{g},\theta\mid\tilde{M})$ $=\sum_{i=1}^{g}\Bigg{[}m_{i0}\log\left(\left(2(1-\pi_{i})^{-\theta}-1\right)^{-\frac{1}{\theta}}\right)$
@@
$\ell_{0}(\pi_{0},\theta\mid\tilde{M})$ $=\sum_{i=1}^{g}\Bigg{[}m_{i0}\log\left(\left(2(1-\pi_{0})^{-\theta}-1\right)^{-\frac{1}{\theta}}\right)$
@@ 2.3.1 Likelihood Ratio Test ($T_{LR}$)
$T_{LR}=2(\ell(\hat{\pi}_{1},\dots,\hat{\pi}_{g},\hat{\theta}\mid\tilde{M})-\ell_{0}(\pi_{0},\hat{\theta}\mid\tilde{M})).$
Under $H_{0}$, $T_{LR}$ follows asymptotically a chi-square distribution with $g-1$ degrees of freedom (Wilks,, 1938).
@@ 2.3.2 Score Test ($T_{S}$)
The expression of $I_{ii}$, $I_{i(g+1)}$, and $I_{(g+1)(g+1)}$ ($i=1,2,...,g$) can be found in the Appendix. The score test statistic $T_{S}$ can be further simplified as:
@@
$d_{i}=\;-\;m_{\textrm{i1}}\,\frac{\displaystyle\left(\,4/\omega(\pi_{i})^{\tfrac{1}{\theta}+1}\,(1-\pi_{i})^{\theta+1}\;-\;2\right)}{\displaystyle 2\,\pi_{i}\;+\;{2}/{\omega(\pi_{i})^{\tfrac{1}{\theta}}}\;-\;2}\;-\;m_{\textrm{i2}}\,\frac{\displaystyle\left(\,{2}/{\omega(\pi_{i})^{\tfrac{1}{\theta}+1}\,(1-\pi_{i})^{\theta+1}}\;-\;2\right)}{\displaystyle 2\,\pi_{i}\;+\;{1}/{\omega(\pi_{i})^{\tfrac{1}{\theta}}}\;-\;1}\;-\;\frac{2\,m_{\textrm{i0}}}{2-2\,\pi_{i}\;-\;\bigl{(}1-\pi_{i}\bigr{)}^{\theta+1}}\,,$
with $\text{with }\omega(\pi_{i})\;=\;{2}/{\bigl{(}1-\pi_{i}\bigr{)}^{\theta}}\;-\;1.$
$T_{S}$ is asymptotically distributed as a chi-square distribution with $g-1$ degrees of freedom according to Rao, (1948).
@@ 2.3.3 Wald Test ($T_{W}$)
which follows asymptotically a $\mathcal{X}^{2}$ distribution with $g-1$ degrees of freedom under the alternative hypothesis $H_{a}$ (Wald,, 1943).
@@ Appendix A Fisher Information Matrix
×$\times{\left(3\pi_{i}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-2{\pi_{i}}^{2}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}+1\right)}$
@@
×$\times\Biggl{[}\pi_{i}{(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}})}^{1/\theta}-{(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}})}^{1/\theta}$

### v2

# Testing the Homogeneity of Proportions for Correlated Bilateral Data via the Clayton Copula

@@ 2.1 Copula Model
[…]
@@ 2.2 Likelihood Inference
[…]
@@
[…]
@@
[…]
@@ 2.3.1 Likelihood Ratio Test ($T_{LR}$)
$).T_{LR}=2(\ell(\hat{\pi}_{1},\dots,\hat{\pi}_{g},\hat{\theta}\mid\tilde{M})-\ell_{0}(\pi_{0},\hat{\theta}\mid\tilde{M}$)).
Under $H_{0}$, $T_{LR}$ follows asymptotically a chi-square distribution with g1g-1 degrees of freedom (Wilks,, 1938).
@@ 2.3.2 Score Test ($T_{S}$)
The expression of $I_{ii}$, $I_{i(g+1)}$, and $I_{(g+1)(g+1)}$ (i=1,2,,gi=1,2,...,g) can be found in the Appendix. The score test statistic $T_{S}$ can be further simplified as:
@@
[…]
$\text{with }\omega(\pi_{i})\;=\;{2}/{\bigl(1-\pi_{i}\bigr)^{\theta}}\;-\;1.$
$T_{S}$ is asymptotically distributed as a chi-square distribution with g1g-1 degrees of freedom according to Rao, (1948).
@@ 2.3.3 Wald Test ($T_{W}$)
which follows asymptotically a $\mathcal{X}^{2}$ distribution with g1g-1 degrees of freedom under the alternative hypothesis $H_{a}$ (Wald,, 1943).
@@ Appendix A Fisher Information Matrix
$\times{\left(3\pi_{i}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-2{\pi_{i}}^{2}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}+1\right)}$
@@
$\times\Biggl[\pi_{i}{(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}})}^{1/\theta}-{(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}})}^{1/\theta}$